I am having problems completing the square x^2-8x=10. Below is the steps that I have taken:
x^2-8x+16=10+16
x^2-8x =26
(x-4)^2=26
This is where I am completely lost.

Found 2 solutions by fractalier, stanbon:
Answer by fractalier(6550)

From x^2 - 8x = 10, you are right if you have
x^2 - 8x + 16 = 10 + 16
x^2 - 8x + 16 = 26
(x - 4)^2 = 26
Now take the square root of both sides and solve for x, giving you
x - 4 = ± sqrt(26)
x = 4 ± sqrt(26)
and we're done...

Answer by stanbon(75887)

I am having problems completing the square x^2-8x=10. Below is the steps that I have taken:
x^2-8x=10
x^2-8x+16=10+16
You have to notice now that the left side is a perfect square, so factor it, as follows:
(x-4)^2=26
Now, take the square root of both sides to get:
x-4=+-sqrt26
x=4+sqrt(26 or x=4-sqrt26
Check these solutions in the original equation.
Cheers,
Stan H.